The test scores of 40 students are summarized in the frequency distribution below. Find the standard deviation. Score Students 50-59 5 60-69 8 70-79 6 80-89 13 90-99 8


1
Expert's answer
2021-10-12T11:14:57-0400


Mean "\bar{X} = \frac{\sum f_iX_i}{N} = \frac{3090}{40}=77.25"

Standard deviation "= \sqrt{\frac{\sum f_iX_i^2}{N} - (\bar{X})^2}"

"= \sqrt{ \frac{245700}{40} -(77.25)^2} \\

= \sqrt{6142.5 -5967.56} \\

= \sqrt{174.94} \\

= 13.22"